If have a 50 lb. bag of dry powder that is to be mixed with 6.5 quarts of water, but I only need to mix a very small amount of the product.
If I want to use only 1/2 lb. of powder, how much water would I use?

Answer by ptaylor(2198)

Let x=amount of water needed
Now we can set up a simple ratio:
50 is to 6.5 as 0.5 is to x and we write this as follows:
50/6.5=0.5/x multiply each term by 6.5x (or simply cross-multiply)
50x=3.25
x= 0.065 qts per 1/2lb----amount of water needed
CK
(100)*(1/2)gal=50gal
100*0.065=6.5 qts
Hope this helps---ptaylor